# Admissible covers and stable maps

### v1

@@ 1.1. Results
Gromov–Witten cycles can be viewed as associated to the loci of genus $g$ curves admitting a degree $d$ stable map [^1] A map is stable, if it is not fixed by automorphisms of curves. to a genus $h$ curve,
@@
The first two results correspond to degree $d=1$ maps to the moving elliptic curve and degree $d=2$ maps to $\mathbb{P}^{1}$ with two relative points, respectively. In Appendix A written by Johannes Schmitt, we give a low-genus computational verification of the formula in the case of degree $d=1$ and $d=2$ maps to $\mathbb{P}^{1}$ with two relative points, using admcycles [DSvZ21]. The third result follows from taking the minimal power of $z$ in the formula associated to maps to $\mathbb{P}^{1}$ with one relative point, Theorem 6.1.
@@ 2.5. Star-shaped graphs
Figure 1. Star-shaped graph
@@ 2.7. Hodge classes
which admit an expression in terms of $\lambda$- and $\psi$-classes, obtained by analysing $e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})$.
@@ 2.8. Wall-crossing formula
where the sum is taken over all ways of attaching contracted components to a cover. In particular, these components may be attached either away from or at the ramification locus. In this perspective, $I$-functions are precisely the (virtual) classes associated to moduli spaces of contracted components of curves.
@@ 3.1. Idea of the proof
The result follows from the localisation on the master space associated to $\epsilon$-unramification, defined in full generality in [Nes24] and considered in a slightly different context in [Nes22]. [^3] The set-up of [Nes22] did not include markings in the source curves, and only relative insertions were permitted. In fact, Theorem 2.7 is a cycle-valued upgrade of the numerical wall-crossing discussed in [Nes24]. The proof is quite similar with the difference that we project our classes to ${\overline{M}}_{g,n}$ instead of a point. We establish a wall-crossing formula for one wall in Section 3.6. Theorem 2.7 is the result of applying (3.5) to all walls on the way from stable maps to admissible covers.
@@ 3.2. Stability
Let $(C,\mathbf{p})$ and $(X,\mathbf{x})$ be two marked nodal curves. Consider an admissible map,
@@
be the moduli space of $\epsilon$-unramified maps of degree $d$ from possibly disconnected curves of genus $g$ with $n$ marked points, and with ramifications over $m$ marked points on the target specified a vector of partitions $\underline{\mu}:=(\mu^{1},\dots,\mu^{m})$.
@@ 3.3. Master space
On the other hand, as explained in [Nes24, Section 6.7], elements of $F_{\Gamma}$ are maps to curves with rational tails, such that the shape of a map (i.e., how the genus and ramifications are redistributed) is determined by the graph $\Gamma$. The union is taken over all star-shaped graphs $\Gamma$ of weight $d_{0}$, that is
@@
Figure 2. Star-shaped graph with leaves at the root
Let $F^{\mathrm{ord}}_{\Gamma}$ be the space associated to $F_{\Gamma}$ by putting a standard order (i.e., the order which is in agreement with genus labels of $\Gamma$) on rational tails and on the ramification points over nodes. Then the space $F^{\mathrm{ord}}_{\Gamma}$ is a ${\mathbb{Z}}_{k}$-gerbe over the following product,
@@ 3.4. Localisation on the master space
given by applying stabilisation to all rational tails of target curves and composing maps with the stabilisation morphism. We could go even further down to ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$, but pullbacks of $\psi$-classes from ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$ and relative $\psi$-classes on $M{\overline{M}}^{\epsilon_{0}}_{g,n}({\mathcal{X}},\underline{\mu})$ do not agree in general, due to ramifications away from the marked points on the target. These are retained in ${\overline{M}}_{g,n}({\mathcal{X}},\underline{\mu})$ but forgotten in ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$. Nevertheless, they do agree in the case of admissible covers, which will allow us to consider the forgetful morphism to ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$ in the final chamber of the wall-crossing.
@@
Since $M{\overline{M}}^{\epsilon_{0}}_{g,n}({\mathcal{X}},\underline{\mu})$ is proper, the left-hand side of the relation above does not involve negative powers of $z$. After taking the residue at $z=0$, i.e., the coefficients of $\frac{1}{z}$, we therefore obtain
@@
it remains to analyse the residue of the wall-crossing components $F_{\Gamma}$. Note that since we taking the residue, a precise form of lines bundles $L_{-}$ and $L_{+}$ is not required.
@@ 3.5. Residue
(3.2) $\mathrm{Res}_{z}\pi_{*}\left(\frac{[F_{\Gamma}]^{\mathrm{vir}}\cap A}{e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})}\right)=\\ \mathrm{Res}_{z}\frac{1}{\operatorname{Aut}(\Gamma)}\cdot\frac{1}{\operatorname{Aut}(\underline{\eta})}\cdot\mathrm{gl}_{\Gamma*}\left(\pi_{*}\left(\frac{[\widetilde{M}^{\epsilon_{+}}_{g_{0},n_{0}}(X,\underline{\mu},\underline{\eta}]^{\mathrm{vir}}\cap A_{\Gamma})}{-z-\sum^{\infty}_{i=0}D_{i}}\right)\boxtimes\prod^{k}_{i=1}\mathsf{I}_{g_{i},n_{i},\eta^{i}}(z-\tilde{\psi_{i}})\right),$
@@
where $\tilde{\operatorname{ev}}^{\prime}$ is the evaluation map associated to branching points with the ramification profile $\underline{\eta}$; the class $\gamma_{\Gamma}$ is defined as in (2.3). The factor $\operatorname{Aut}(\Gamma)$ is due to the ordering of rational tails and ramification points in $F_{\Gamma}$. Lastly, the classes $\tilde{\psi}_{i}$ in the expression above are associated to nodes of target curves. On moduli spaces of maps with relative points they admit an identification with $\psi$-classes on the source via the formula involving a part of the partition given in the end of Section 2.5.
@@
(3.3) $\sum_{w(\Gamma)=d_{0}}\mathrm{Res}_{z}\pi_{*}\left(\frac{[F_{\Gamma}]^{\mathrm{vir}}\cap A_{\Gamma}}{e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})}\right)=\sum_{w(\Gamma_{\mathrm{ord}})=d_{0}}\sum^{k-1}_{r=0}\sum_{\underline{b}}\frac{(-1)^{r}}{r!(k-r)!}\cdot\frac{1}{\operatorname{Aut}(\underline{\eta})}\cdot\\ \mathrm{gl}_{\Gamma_{\mathrm{ord}}*}\bigg{(}\pi_{*}\left([{\overline{M}}^{\epsilon_{+}}_{g_{0},n_{0}}({\mathcal{X}},\underline{\mu},\underline{\eta})]^{\mathrm{vir}}\cap A_{\Gamma_{\mathrm{ord}}}\right)\boxtimes\prod^{k}_{i=1}[\mathsf{I}_{g_{i},n_{i},\eta^{i}}(z-\tilde{\psi_{i}})]_{b_{\ell}}\bigg{)},$
@@
Remark 3.3.
Strictly speaking, to obtain (3.3) from (3.2), it is also necessary to consider the wall-crossing for $(h,m)=(0,1)$, which is presented in the proof of Theorem 6.1. The reason is that the contributions of $D_{i}$ are expressible in terms of polar parts of $I$-functions, as explained in [Zho22, Lemma 7.2.1].
@@ 4.2. Genus three
where $[\alpha_{1},\dots,\alpha_{k}]_{g_{0},g_{1},\ldots,g_{k}}=\mathrm{gl}_{*}(\alpha_{0}\boxtimes\alpha_{1}\boxtimes\dots\boxtimes\alpha_{k}$). Over the locus of compact-type curves $M^{\mathrm{ct}}_{3}\subset{\overline{M}}_{3}$, using the admcycles [DSvZ21], the expression above yields
@@ 4.3. Genus four
Over the locus of compact-type curves $M^{\mathrm{ct}}_{4}\subset{\overline{M}}_{4}$, using the admcycles package [DSvZ21], the expression above yields
@@ 5.3. Genus three
$\mathsf{I}_{2,(2)}(z)=2z^{-1}\cdot\frac{z^{2}-\lambda_{1}z+\lambda_{2}}{\frac{z}{2}-\psi_{1}}=4\mathbbm{1}+\mathrm{O}(z^{-1}),$
@@
As explained in Appendix A, using the admcycles [DSvZ21], Corollary 5.1 was verified for $g\leq 3$. For computations, it is important to include all vertices of graphs, even those corresponding to $\mathsf{I}_{0,(2)}(z)$, because of the insertion of $\psi$-classes on root components. [^5] A $\psi$-class and the pullback of $\psi$-class by a forgetful map differ by a boundary divisor. The same applies to conjugate points - we can forget a conjugate point which is not glued to any other components only after inserting $\psi$-classes.
@@ 6.1. The polar part of $I$-function
Note that on the left from the wall $\epsilon_{0}=2g-2+d+\ell(\mu)+n$, $\epsilon$-unramification gives an empty moduli space by Definition 3.1. Moreover, the master space for the wall $\epsilon_{0}$ takes a different and simpler form. It is defined as follows.
Let $\mathfrak{M}_{g,n}(\mathbb{P}^{1}/\infty,\mu)^{\sim}$ the moduli stack of maps to $\mathbb{P}^{1}$ with ramification $\mu$ over $\infty\in\mathbb{P}^{1}$ up to $\operatorname{Aut}(\mathbb{P}^{1},\infty)$, such that the automorphisms of maps to the fixed $\mathbb{P}^{1}$ are finite (i.e., we require that the automorphisms of maps to $\mathbb{P}^{1}$ are finite, but they might be infinite if automorphisms of the target $\operatorname{Aut}(\mathbb{P}^{1},\infty)$ are taken into account). Let $L$ be the line bundle given by the cotangent space at $\infty\in\mathbb{P}^{1}$. Consider the projective bundle,
@@
More explicitly, the points of this space are triples $(f,v_{1},v_{2})$, where $f\in\mathfrak{M}_{g,n}(\mathbb{P}^{1}/\infty,\mu)^{\sim}$, and $v_{1}\in H^{0}(C,L\otimes N)$ and $v_{2}\in H^{0}(C,N$) are sections with no common zeros on the source curve $C$ together with a line bundle $N$. We define
@@ 6.2. ELSV formula
which is the ELSV formula. Theorem 6.1 can be viewed as its cycle-valued refinement.
@@ 7.4. Fulton–MacPherson invariants
Once defined correctly, a moduli space of admissible covers admits an étale map to a Fulton–MacPherson space [FM94] of $X$ by associating branching points on the target to an admissible cover,
(7.1) $\overline{H}_{\underline{\eta}}(X)\rightarrow{FM}_{k}(X),$
@@
$\operatorname{ev}\colon FM_{k}(X)\rightarrow X^{k}.$
We define invariants associated to $FM_{k}(X)$,
$\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{m}}(\gamma_{m})\rangle^{\mathsf{FM}}:=\int_{[FM_{k}(X)]}\operatorname{ev}^{*}(\gamma_{1}\boxtimes\dots\boxtimes\gamma_{n})\cdot\prod\psi^{k_{i}}_{i}.$
@@
(7.2) $\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{m}}(\gamma_{m})\rangle^{\mathsf{H}}_{\underline{\eta}}=\mathrm{Hur}(\underline{\eta})\cdot\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{m}}(\gamma_{m})\rangle^{\mathsf{FM}}.$
@@
$\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{m}}(\gamma_{m})\cdot\mathbbm{1}\rangle^{\mathsf{FM}}=\sum^{n}_{i=1}\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{i}-1}(\gamma_{i})\dots\tau_{k_{m}}(\gamma_{m})\rangle^{\mathsf{FM}}.$

### v2

@@ 1.1. Results
Gromov–Witten cycles can be viewed as associated to the loci of genus $g$ curves admitting a degree $d$ stable map [^1] A map is stable, if it is fixed by at most finitely many automorphisms of target and source curves. to a genus $h$ curve,
@@
The first two results correspond to degree $d=1$ maps to the moving elliptic curve and degree $d=2$ maps to $\mathbb{P}^{1}$ with two relative points, respectively. In Appendix A written by Johannes Schmitt, we give a low-genus computational verification of the formula in the case of degree $d=1$ and $d=2$ maps to $\mathbb{P}^{1}$ with two relative points, using the admcycles package [DSvZ21]. The third result follows from taking the minimal power of $z$ in the formula associated to maps to $\mathbb{P}^{1}$ with one relative point, Theorem 6.1.
@@ 2.5. Star-shaped graphs
Figure 1. Star-shaped graph.
@@ 2.7. Hodge classes
which admits an expression in terms of $\lambda$- and $\psi$-classes, obtained by analysing $e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})$.
@@ 2.8. Wall-crossing formula
where the sum is taken over all ways of attaching contracted components to a cover. In particular, these components may be attached either away from or at the ramification locus. In this perspective, $I$-functions are precisely the (virtual) classes associated to moduli spaces of contracted components of curves. The case $(h,m)=(0,1)$ admits a different wall-crossing formula, which is considered in Section 6.
@@ 3.1. Idea of the proof
The result follows from the localisation on the master space associated to $\epsilon$-unramification, defined in full generality in [Nes24] and considered in a slightly different context in [Nes22]. [^3] The set-up of [Nes22] did not include markings in the source curves, and only relative insertions were permitted. In fact, Theorem 2.7 is a cycle-valued upgrade of the numerical wall-crossing discussed in [Nes24]. The proof is quite similar, the difference is that we project our classes to ${\overline{M}}_{g,n}$ instead of a point. We establish a wall-crossing formula for one wall in Section 3.6. Theorem 2.7 is obtained by applying (3.5) to all walls along the path from stable maps to admissible covers.
@@ 3.2. Stability
Let $(C,\mathbf{p})$ and $(X,\mathbf{x})$ be two marked nodal curves. Consider a map,
@@
be the moduli space of $\epsilon$-unramified maps of degree $d$ from possibly disconnected curves of genus $g$ with $n$ marked points, and with ramifications over $m$ marked points on the target specified a vector of partitions $\underline{\mu}:=(\mu^{1},\dots,\mu^{m})$. These are proper spaces by [Nes24, Theorem 3.15]; see also [Nes22], where the case of a moving curve without markings on source curves was considered.
@@ 3.3. Master space
On the other hand, as explained in [Nes24, Section 6.7], elements of $F_{\Gamma}$ are maps to curves with rational tails of degree $d_{0}$ with respect to the branching divisor, such that the shape of a map (i.e., how the genus and ramifications are redistributed) is determined by the graph $\Gamma$; over the distinguished degree-$d_{0}$ rational tails specified by non-root vertices of the graph, the map belongs to $V^{{\mathbb{C}}^{*}}_{g_{i},n_{i},\eta^{i}}$. The union is taken over all star-shaped graphs $\Gamma$ of weight $d_{0}$, that is
@@
Figure 2. Star-shaped graph with leaves at the root.
Let $F^{\mathrm{ord}}_{\Gamma}$ be the space associated to $F_{\Gamma}$ by putting a standard order (i.e., the order which is in agreement with genus labels of $\Gamma$) on rational tails of degree $d_{0}$ and on the ramification points over nodes. Then the space $F^{\mathrm{ord}}_{\Gamma}$ is a ${\mathbb{Z}}_{k}$-gerbe over the following product,
@@ 3.4. Localisation on the master space
given by applying stabilisation to all rational tails of target curves, composing maps with the stabilisation morphism, and stabilising the source curves. We could go even further down to ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$, but pullbacks of $\psi$-classes from ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$ and relative $\psi$-classes on $M{\overline{M}}^{\epsilon_{0}}_{g,n}({\mathcal{X}},\underline{\mu})$ do not agree in general due to ramifications away from the marked points on the target. These are retained in ${\overline{M}}_{g,n}({\mathcal{X}},\underline{\mu})$ but forgotten in ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$. Nevertheless, they do agree in the case of admissible covers, which will allow us to consider the forgetful morphism to ${\overline{M}}_{g,n+\sum\ell(\mu^{i})}$ in the final chamber of the wall-crossing.
@@
Since $M{\overline{M}}^{\epsilon_{0}}_{g,n}({\mathcal{X}},\underline{\mu})$ is proper by [Nes24, Theorem 6.13], the left-hand side of the relation above does not involve negative powers of $z$. After taking the residue at $z=0$, i.e., the coefficients of $\frac{1}{z}$, we therefore obtain
@@
it remains to analyse the residue of the wall-crossing components $F_{\Gamma}$; note that since we are taking the residue, a precise form of lines bundles $L_{-}$ and $L_{+}$ is not required.
@@ 3.5. Residue
(3.2) $\mathrm{Res}_{z}\pi_{*}\left(\frac{[F_{\Gamma}]^{\mathrm{vir}}\cap A}{e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})}\right)=\\ \mathrm{Res}_{z}\frac{1}{\operatorname{Aut}(\Gamma)}\cdot\frac{1}{\operatorname{Aut}(\underline{\eta})}\cdot\mathrm{gl}_{\Gamma*}\left(\pi_{*}\left(\frac{[\widetilde{M}^{\epsilon_{+}}_{g_{0},n_{0}}(X,\underline{\mu},\underline{\eta})]^{\mathrm{vir}}\cap A_{\Gamma})}{-z-\sum^{\infty}_{i=0}D_{i}}\right)\boxtimes\prod^{k}_{i=1}\mathsf{I}_{g_{i},n_{i},\eta^{i}}(z-\tilde{\psi_{i}})\right),$
@@
where $\tilde{\operatorname{ev}}^{\prime}$ is the evaluation map associated to the branching points with the ramification profile $\underline{\eta}$; the class $\gamma_{\Gamma}$ is defined as in (2.3). The factor $\operatorname{Aut}(\Gamma)$ is due to the ordering of rational tails and ramification points in $F_{\Gamma}$. Lastly, the classes $\tilde{\psi}_{i}$ in the expression above are associated to nodes of target curves. On moduli spaces of maps with relative points they admit an identification with $\psi$-classes on the source via the formula involving a part of the partition given in the end of Section 2.5.
@@
(3.3) $\sum_{w(\Gamma)=d_{0}}\mathrm{Res}_{z}\pi_{*}\left(\frac{[F_{\Gamma}]^{\mathrm{vir}}\cap A}{e_{{\mathbb{C}}^{*}}(N^{\mathrm{vir}})}\right)=\sum_{w(\Gamma_{\mathrm{ord}})=d_{0}}\sum^{k-1}_{r=0}\sum_{\underline{b}}\frac{(-1)^{r}}{r!(k-r)!}\cdot\frac{1}{\operatorname{Aut}(\underline{\eta})}\cdot\\ \mathrm{gl}_{\Gamma_{\mathrm{ord}}*}\bigg{(}\pi_{*}\left([{\overline{M}}^{\epsilon_{+}}_{g_{0},n_{0}}({\mathcal{X}},\underline{\mu},\underline{\eta})]^{\mathrm{vir}}\cap A_{\Gamma_{\mathrm{ord}}}\right)\boxtimes\prod^{k}_{i=1}[\mathsf{I}_{g_{i},n_{i},\eta^{i}}(z-\tilde{\psi_{i}})]_{b_{\ell}}\bigg{)},$
@@
Remark 3.3.
To obtain (3.3) from (3.2), it is also necessary to consider the wall-crossing for $(h,m)=(0,1)$, which is presented in the proof of Theorem 6.1. The reason is that the contributions of $D_{i}$ are expressible in terms of polar parts of $I$-functions, as explained in [Zho22, Lemma 7.2.1].
@@ 4.2. Genus three
where $[\alpha_{1},\dots,\alpha_{k}]_{g_{0},g_{1},\ldots,g_{k}}=\mathrm{gl}_{*}(\alpha_{0}\boxtimes\alpha_{1}\boxtimes\dots\boxtimes\alpha_{k}$). Over the locus of compact-type curves $M^{\mathrm{ct}}_{3}\subset{\overline{M}}_{3}$, using admcycles [DSvZ21], the expression above yields
@@ 4.3. Genus four
Over the locus of compact-type curves $M^{\mathrm{ct}}_{4}\subset{\overline{M}}_{4}$, using admcycles [DSvZ21], the expression above yields
@@ 5.3. Genus three
$\mathsf{I}_{2,(2)}(z)=2z^{-1}\cdot\frac{z^{2}-\lambda_{1}z+\lambda_{2}}{\frac{z}{2}-\psi_{1}}=4\mathbbm{1}+\mathrm{O}(z^{-1}).$
@@
As explained in Appendix A, using admcycles [DSvZ21], Corollary 5.1 was verified for $g\leq 3$. For computations, it is important to include all vertices of graphs, even those corresponding to $\mathsf{I}_{0,(2)}(z)$, because of the insertion of $\psi$-classes on root components. [^5] A $\psi$-class and the pullback of $\psi$-class by a forgetful map differ by a boundary divisor. The same applies to conjugate points - we can forget a conjugate point which is not glued to any other components only after inserting $\psi$-classes.
@@ 6.1. The polar part of $I$-function
Note that on the left from the wall $\epsilon_{0}=2g-2+d+\ell(\mu)+n$, the conditions of $\epsilon$-unramification give an empty moduli space by Definition 3.1. Moreover, the master space for the wall $\epsilon_{0}$ takes a different and simpler form. It is defined as follows.
Let $\mathfrak{M}_{g,n}(\mathbb{P}^{1}/\infty,\mu)^{\sim}$ the moduli stack of maps to $\mathbb{P}^{1}$ with the ramification profile $\mu$ over $\infty\in\mathbb{P}^{1}$ up to $\operatorname{Aut}(\mathbb{P}^{1},\infty)$, such that the automorphisms of maps to the fixed $\mathbb{P}^{1}$ are finite (i.e., we require that the automorphisms of maps to $\mathbb{P}^{1}$ are finite, but they might be infinite if automorphisms of the target $\operatorname{Aut}(\mathbb{P}^{1},\infty)$ are taken into account). Let $L$ be the line bundle given by the cotangent space at $\infty\in\mathbb{P}^{1}$. Consider the projective bundle,
@@
The $B$-points of this space are tuples $(f,N,v_{1},v_{2})$, where $f\in\mathfrak{M}_{g,n}(\mathbb{P}^{1}/\infty,\mu)^{\sim}(B)$, $N$ is a line bundle on the base scheme $B$, and $v_{1}\in H^{0}(B,L_{B}\otimes N)$ and $v_{2}\in H^{0}(B,N$) are sections with no common zeros. We define
@@ 6.2. ELSV formula
which is the ELSV formula, proved in [ELSV01, GV03]. Theorem 6.1 can be viewed as its cycle-valued refinement.
@@ 7.4. Fulton–MacPherson invariants
A moduli space of admissible covers admits a finite map to a Fulton–MacPherson space [FM94] of $X$ by associating branching points on the target to an admissible cover,
(7.1) $\overline{H}_{\underline{\eta}}(X)\rightarrow{FM}_{n}(X),$
@@
$\operatorname{ev}\colon FM_{n}(X)\rightarrow X^{n}.$
We define invariants associated to $FM_{n}(X)$,
$\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{n}}(\gamma_{n})\rangle^{\mathsf{FM}}:=\int_{[FM_{n}(X)]}\operatorname{ev}^{*}(\gamma_{1}\boxtimes\dots\boxtimes\gamma_{n})\cdot\prod^{n}_{i=1}\psi^{k_{i}}_{i}.$
@@
(7.2) $\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{n}}(\gamma_{n})\rangle^{\mathsf{H}}_{\underline{\eta}}=\mathrm{Hur}(\underline{\eta})\cdot\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{n}}(\gamma_{n})\rangle^{\mathsf{FM}}.$
@@
$\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{n}}(\gamma_{n})\cdot\mathbbm{1}\rangle^{\mathsf{FM}}=\sum^{n}_{i=1}\langle\tau_{k_{1}}(\gamma_{1})\dots\tau_{k_{i}-1}(\gamma_{i})\dots\tau_{k_{n}}(\gamma_{n})\rangle^{\mathsf{FM}}.$
Proof.
The result follows by applying forgetful maps between Fulton–MacPherson spaces, and using the relation between a $\psi$-class and the pullback of a $\psi$-class. ∎